OpenStudy (anonymous):

Kyle was trying to improve his time in the 200 meter dash. He started at 30.5 seconds, then ran 29.9 seconds, 29.4 seconds and 29 seconds. If he continues practicing, what is a reasonable prediction for his next time?

OpenStudy (anonymous):

30 seconds 28.7 seconds 26 seconds 28 seconds

OpenStudy (anonymous):

Find the difference from one time to the next. See if there is a pattern.

OpenStudy (anonymous):

30.5 -29.9=.6 29.9-29.4=.5 29.4-29=.4 hence next time it will be 29 - 0.3 =28.7